# IMPROVING AUDIO GENERATION WITH VISUAL ENHANCED CAPTION

### v1

@@ 5 Conclusion
This paper presents a large audio-caption dataset to address the challenge of prompt following in audio generation systems. Specifically, we propose Sound-VECaps, a large-scale dataset comprising over 1.66 million audio clips with captions enriched by visual feature guidance. To evaluate its effectiveness, we developed new benchmarks using AudioCaps testing sets with improved and enriched captions.
Our AudioLDM model with Sound-VECaps achieves state-of-the-art performance and outperforms baseline models. In addition, our system further improves by a large margin when generating complex and enhanced prompts, reaching a FAD score of . The proposed datasets demonstrate substantial improvements in audio retrieval and temporal feature identification. Nevertheless, the results between AudioCaps and AudioCaps-Enhanced testing sets highlight the limitations of previous benchmarks that rely on simpler prompts and emphasize the value and potential of the proposed benchmark in advancing audio-language model performance. For both the Sound-VECaps and AudioCaps-Enhanced testing set, we developed both audio-only versions and captions with full features for different purposes and tasks. Although the improvements from the dataset are confined to specific audio-language tasks, we hope these datasets can offer advancements in audio generation systems and support the community in making future progress in audio-language tasks.

### v3

@@ 5 Conclusion
This paper presents Sound-VECaps, a large-scale dataset comprising over M audio clips with captions enriched by visual feature guidance, to address the challenge of prompt following in audio generation systems. In addition, a new benchmark using improved and enriched captions is proposed to evaluate audio-language systems on complex and extended prompts.
Experiments show that AudioLDM models trained on Sound-VECaps achieve SoTA performance and outperform baseline models. In addition, our systems further improve by a large margin when taking more detailed captions as prompts, reaching a FAD score of . In addition, Sound-VECaps demonstrate substantial improvements in audio retrieval and temporal feature identification. Nevertheless, the results between AudioCaps and AudioCaps-Enhanced testing sets highlight the limitations of previous benchmarks that rely on simple prompts and emphasize the value and potential of the more detailed prompts in advancing the performance of audio-language models. For both the Sound-VECaps and AudioCaps-Enhanced testing set, we developed audio-only versions and full versions with captions that include and exclude visual-only content for different purposes and tasks. Although the improvements from the dataset are confined to specific audio-language tasks, we hope these datasets can offer advancements in audio generation systems and support the community in making future progress in audio-language tasks.
6 ACKNOWLEDGMENT
This research was partly supported by a research scholarship from the China Scholarship Council (CSC), an internship from ByteDnace, and supported by the Engineering and Physical Sciences Research Council (EPSRC) under Grant EP/T019751/1 “AI for Sound”.